### 21 Dec 2023 v1

# Blown-up boundaries associated with ample cones of K3 surfaces

Abstract.
In this paper, we apply the Bestvina-Mess type formula for relatively hyperbolic groups, which is established by Tomohiro Fukaya, to automorphism groups of K3 surfaces, and we show that the virtual cohomological dimension of automorphism groups of K3 surfaces determined by the covering dimension of the blown-up boundaries associated with their ample cones.
Key words and phrases:
K3 surfaces, Automorphism groups of K3 surfaces, Cohomological dimension, Blown-up boundaries of ample cones, Sphere packings
2020 Mathematics Subject Classification:
Primary 14J28; Secondary 14J50, 14J27
Department of Mathematics, Tokyo Institute of Technology, Tokyo 152-8551, Japan.
e-mail: takatsu.t.aa@m.titech.ac.jp
1. Introduction
In this paper, we discuss the classification problem of automorphism groups of K3 surfaces over $\mathbb{C}$.
There are two approaches to automorphisms of K3 surfaces. The first one is the classification under some finiteness conditions, such as the classification of finite subgroups, and that of possible finite order of automorphisms. For example, a well-known and important result in this approach is Mukai’s result [16] Theorem (0.3), which says that symplectic finite subgroups of the automorphism groups of K3 surfaces are isomorphic to subgroups of the Mathieu group $M_{23}$, one of the sporadic simple groups.
The other approach is the application of Borcherds’ lattice theory (often called Borcherds’ method) to compute a generating set of infinite automorphism groups. The problem of finding a generating set of the automorphism groups of Kummer surfaces was first proposed by Klein. This problem took more than 100 years to be solved by Kondo [12] using Borcherds’ method. Although this method is very useful and, in fact, the automorphism groups of some specific K3 surfaces have been calculated, this method is not applicable to the general case.
Shigeru Mukai recently proposed at the occasion of the Algebra Symposium in Tokyo in 2018 the following conjecture on discrete group theory and automorphism groups, from which one may hopefully find a general framework for dealing with infinite automorphism groups of elliptic K3 surfaces.
Conjecture 1.
Let $X$ be an elliptic K3, Enriques or Coble surface, and let $\operatorname{Aut}(X)$ be the automorphism group of X. Then, the virtual cohomological dimension $\operatorname{vcd}(\operatorname{Aut}(X))$ of $\operatorname{Aut}(X)$ is equal to the maximum of the Mordell-Weil rank induced by all elliptic surface structures on $X$.
Remark 1.1.
The virtual cohomological dimension of $\operatorname{Aut}(X)$ is easily seen to be larger than the rank of the Mordell-Weil group of any elliptic fibration. Hence the main point of Mukai’s conjecture lies in finding the upper-bound of $\operatorname{vcd}(\operatorname{Aut}(X))$ as stated therein.
This conjecture could be a turning point from the conventional study of automorphism groups of elliptic K3, Enriques and Coble surfaces by computing their generators to that by computing their cohomological invariants. Our main result is a formula which determines the virtual cohomological dimension of the automorphism group of a K3 surface by the covering dimension of the blown-up boundary associated with the ample cone of the K3 surface.
Theorem 1.
If $\operatorname{Aut}_{s}(X)$ is non-elementary, then
$\operatorname{vcd}(\operatorname{Aut}(X))=\dim\partial A_{X}^{bl}+1,$
where $\partial A_{X}^{bl}$ is the blown-up boundary associated with the ample cone of $X$ (see Section 5.2 for the detailed discussion).
To show this result, we focus on the natural action of automorphism groups on hyperbolic spaces, which come from the action on the cohomology of degree 2. Then we can consider the symplectic automorphism group of a K3 surface as a discrete subgroup of Möbius transformations of a hyperbolic space. Moreover, we can show the following theorem:
Theorem 2.
Let $X$ be a K3 surface, and let $\operatorname{Aut}_{s}(X)$ be the symplectic automorphism group of $X$. Then $\operatorname{Aut}_{s}(X)$ is geometrically finite.
In [5], Bestvina and Mess showed that the cohomological dimension of a torsion free hyperbolic group is determined by the covering dimension of its Gromov boundary. Recently, Tomohiro Fukaya [10] generalized the formula for relatively hyperbolic groups satisfying a certain condition. Geometrically finite groups are typical examples of relative hyperbolic groups. We can immediately obtain Theorem 1 by applying Fukaya’s result [10, Corollary B].
On the other hand, the paper of Fukaya is still unpublished, and his theorem and his proof are stated using geometric group theory. For readability to algebraic geometers, this paper also contains the result and the proof of Fukaya, which are simplified and modified to fit in with our situation, in Section 5.2, using the geometry of ample cones of K3 surfaces and hyperbolic geometry.
Remark 1.2.
After the author finished the first draft of this paper, Tomohiro Fukaya informed the author that Theorem 2 has been independently obtained and proved by Kohei Kikuta. He showed geometrical finiteness for the automorphism groups of irreducible symplectic varieties over a field of characteristic zero, and his proof is different in some points from our proof given in this paper. For more details, the reader is advised to refer to his forthcoming paper.
An apllication of Theorem 1 is the case that the ample cones of K3 surfaces are sphere packings.
Theorem 3.
Let $X$ be an elliptic K3 surface. If $\partial A_{X}$ is a connected sphere packing, and every elliptic divisor of $X$ corresponds to the tangent point of some two spheres associated with (-2)-curves, then
$\operatorname{vcd}(\operatorname{Aut}(X))=\max\{\operatorname{rk}\operatorname{MW}(f)\}=\rho(X)-3,$
where $f$ runs all elliptic fibrations of $X$ (see Section 5.3 for the detailed discussion).
Baragar [2] constructed K3 surfaces whose ample cones are sphere packings. Then we can compute the virtual cohomological dimension of their automorphism groups. Note that these K3 surfaces provide examples that affirm Conjecture 1, and, to the best of the author’s knowledge, they are the first non-trivial examples in the sense that the verification of the conjecture does not rely on explicit calculation of the group structure of the automorphism groups.
Acknowledgment
First, the author expresses his gratitude to his advisor Fumiharu Kato, for providing the author the interesting topics and for many helpful discussions. He is grateful to Shigeru Mukai for various useful advices on the virtual cohomological dimension of automorphism groups of K3 surfaces. He would also like to thank Takefumi Nosaka, Ryokichi Tanaka, and Kohei Kikuta for valuable comments and careful reading. Finally, we thank Tomohiro Fukaya for many useful comments on hyperbolic geometry and hyperbolic group theory.
2. Lattices and K3 surfaces
In this section, we recall some definitions and facts about lattices and K3 surfaces.
2.1. Lattices
[…]
Let $L$ be a unimodular even lattice, and $S$ be a primitive sublattice of $L$. The orthogonal complement of $S$, denoted by $T$, is also a primitive sublattice of $L$. Let $H$ be a quotient group $L/(S\oplus T)$. We denote by $p_{S}$ (resp. $p_{T}$) the projection map from $A_{S}\oplus A_{T}$ to $A_{S}$ (resp. $A_{T}$).
2.2. K3 surfaces
Theorem 2.1 ([13] Theorem 4.5).
For a K3 surface $X$, the second cohomology group $H^{2}(X,\mathbb{Z})$ with the cup product is a unimodular even lattice with the signature $(3,19)$. Then $H^{2}(X,\mathbb{Z})$ is isomorphic to $U^{\oplus 3}\oplus E_{8}^{\oplus 2}$, where $U$ is the hyperbolic lattice of rank 2.
Definition 2.2.
Let $\omega_{X}$ be a nowhere vanishing holomorphic $2$-form of $X$. The Néron-Severi lattice $NS(X)$ is defined by
$NS(X):=\{x\in H^{2}(X,\mathbb{Z})\hskip 3.00003pt:\hskip 3.00003pt\langle x,\omega_{X}\rangle=0\}=H^{2}(X,\mathbb{Z})\cap H^{1,1}(X,\mathbb{R}).$
The Néron-Severi lattice $NS(X)$ is a lattice by the cup product of signature $(1,\rho(X)-1)$, where $\rho(X)$ is the Picard number of $X$. The transcendental lattice $T_{X}$ of $X$ is defined by $NS(X)^{\bot}$.
Definition 2.3.
Let $X$ be a K3 surface. We define
[…]
and
$\Delta^{+}(X):=\{\delta\in\Delta(X):\delta{\textup{ is effective}}\}.$
For $\delta\in\Delta^{+}(X)$, the reflection $s_{\delta}$ in $H^{1,1}(X,\mathbb{R})$ is defined to be
[…]
The reflection group $W(X)$ is the subgroup of $O(H^{1,1}(X,\mathbb{R}))$ generated by the set of reflections $\{s_{\delta}:\delta\in\Delta(X)\}.$ The cone $P(X)=\{x\in H^{1,1}(X,\mathbb{R}):\langle x,x\rangle>0\}$ has two connected components and the one containing the Kähler class is denoted by $P^{+}(X)$ and called the positive cone.
Definition 2.4.
The group $W(X)$ acts on the space $P^{+}(X)$ in a canonical way. The Kähler cone of $X$ is the one of the fundamental domains is given by
$D(X)=\{x\in P^{+}(X):\langle x,\delta\rangle>0,\forall\delta\in\Delta(X)^{+}\}.$
containing Kähler class of $X$.
Definition 2.5.
Let $X$ be an algebraic K3 surface, The ample cone $A(X)$ is defined to be
$A(X)=\{x\in NS(X)\otimes\mathbb{R}\cap P^{+}(X):\langle x,\delta\rangle>0,\forall\delta\in\Delta(X)^{+}\}.$
Let $f:X\to\mathbb{P}^{1}$ be an elliptic fibration of a K3 surface $X$, and let $j:J\rightarrow\mathbb{P}^{1}$ be the jacobian fibration of $f$. The sections of $j$ generate a group $MW(f)$, which is callded the Mordell-Weil group of the elliptic fibration $f:X\to\mathbb{P}^{1}$. The Mordell-Weil group $MW(f)$ is a finitely generated abelian group. We denote by $r(f)$ the rank of $MW(f)$.
We will use the following formula, which is called Shioda-Tate formula.
Lemma 2.6.
[9, Formula (4.3.2)] Let $f:X\to\mathbb{P}^{1}$ be an elliptic fibration of a K3 surface, and let $F_{v}=f^{-1}(t_{v})\;(1\leq v\leq k)$ be the singular fibres of $f$. We denote by $m_{v}$ the number of irreducible components of $F_{v}$. Then the $r(f)$ is given by the following formula:
$r(f)=\rho(X)-2-\sum_{v=1}^{k}(m_{v}-1).$
3. Hyperbolic geometry
In this section, we recall some basic facts of $n$-dimensional hyperbolic spaces and discrete groups of isometries of $n$-dimensional hyperbolic spaces, which are contained in [18].
Notation
Let $(X,d)$ be a metric space, and let $r$ be a positive real number. Throughout this section, we denote by
$B(a,r)$ the open ball of radius $r$ centered at a point $a$ of $X$,
$C(a,r)$ the closed ball of radius $r$ centered at a point $a$ of $X$,
$S(a,r)$ the sphere of radius $r$ centerd at a point $a$ of $X$,
$N(S,r)$ the $r$-neighborhood of $S$ in $X$,
$I(X)$ the group of isometries of $X$.
We denote by $S^{n}$ the $n$-dimensional unit sphere. A vector $e_{i}$ in $\mathbb{R}^{n}$ is the vector whose $i$-th entry is 1 and the other entries are 00.
In the hyperbolic geometry, it is useful to use some models of hyperbolic spaces for different situations. Therefore, we begin with recalling the hyperboloid model of hyperbolic spaces. In section 3.3 and 3.4 we recall the conformal ball model, and the upper half-plane model of hyperbolic spaces.
3.1. The hyperboloid model
Definition 3.1.
Let $n$ be a non-negative integer. Let $x=(x_{1},\cdots,x_{n},x_{n+1})$ and $y=(y_{1},\cdots,y_{n},y_{n+1})$ be vectors in the $(n+1)$-dimensional real vector space $\mathbb{R}^{n+1}$. The Lorentzian inner product of $x$ and $y$, denoted by $x\circ y$, is defined as follows:
$x\circ y=x_{1}y_{1}+\cdots+x_{n}y_{n}-x_{n+1}y_{n+1}.$
The Lorentzian (n+1)-space is the vector space $\mathbb{R}^{n+1}$ together with the Lorentzian inner product. A real $(n+1)\times(n+1)$ matrix $A$ is said to be Lorentzian if for any $x,y\in\mathbb{R}^{n+1}$, $x\circ y=Ax\circ Ay.$
Definition 3.2.
The hyperboloid model $H^{n}$ of hyperbolic $n$-spaces is defined to be the set $\{x\in\mathbb{R}^{n+1}:x\circ x=-1,\hskip 2.84526ptx_{1}>0\}.$ The hyperbolic distance $d_{H}(x,y)$ between $x$ and $y$ is defined by the equation
$\cosh d_{H}(x,y)=-x\circ y.$
By [18, Theorem 3.2.2], the hyperbolic distance $d_{H}$ is a metric on $H^{n}$. By the definition of $d_{H}$, Lorentz $(n+1)\times(n+1)$ matrices naturally act isometrically on $H^{n}$. We denote the group of isometries by $I(H^{n}).$
Definition 3.3.
Let $H^{n}$ be the hyperboloid model of $n$-dimensional hyperbolic spaces. The set of all Lorentz $(n+1)\times(n+1)$ matrices with matrix multiplication is called the Lorentz group of $(n+1)\times(n+1)$ matrices, denoted by $O(n,1)$. The positive Lorentz group $O^{+}(n,1)$ is defined by
$O^{+}(n,1)=\{A\in O(n,1):AH^{n}=H^{n}\}.$
By [18, Theorem 3.2.3], there is an isomorphism from the positive Lorentzian group $O^{+}(n,1)$ to the group of hyperbolic isometries $I(H^{n})$.
Definition 3.4.
A hyperbolic $m$-plane of $H^{n}$ is the nonempty intersection of $H^{n}$ with an $(m+1)$-dimensional vector subspace of $\mathbb{R}^{n+1}$.
3.2. Möbius transformations
To define the conformal ball model and the upper half-space model of hyperbolic $n$-spaces, we use the hyperboloid model and Möbius transformation. In this section, we recall the definition and some basic facts of Möbius transformation.
Definition 3.5.
Let $a$ be a unit vector in $E^{n}$, and let $t$ be a positive real number. Consider the hyperplane of $E^{n}$ is defined by
$P(a,t)=\{x\in E^{n}:a\cdot x=t\}.$
The reflection $\rho$ of $E^{n}$ in the plane $P(a,t)$ is defined by
$\rho(x)=x+2(t-a\cdot x)a.$
Let $a$ be a point in $E^{n}$, and let $r$ be a positive real number. The sphere of $E^{n}$ of radius $r$ centered at $a$ is defined by
$S(a,r)=\{x\in E^{n}:|x-a|=r\}.$
The reflection (or inversion) $\sigma$ of $E^{n}$ in the sphere $S(a,r)$ is defined by
$\sigma(x)=a+\bigg{(}\frac{r}{|x-a|}\bigg{)}^{2}(x-a).$
Let $\infty$ be a point not in $E^{n}$. We define $\hat{E}^{n}=E^{n}\cup\{\infty\}$ whose topology is given by the one-point compactification of $E^{n}$.
Definition 3.6.
Let $a$ be a unit vector in $E^{n}$, and let $t$ be a positive real number. The extended hyperplane of $\hat{E}^{n}$ is defined by $\hat{P}(a,t)=P(a,t)\cup\{\infty\}.$ Let $\rho$ be the reflection of the hyperplane $P(a,t)$. Extend $\rho$ to a bijection $\hat{\rho}:\hat{E}^{n}\rightarrow\hat{E}^{n}$ by setting $\hat{\rho}(\infty)=\infty.$ In a similar way, the inversion $\sigma$ of an Euclidean sphere $S(a,r)$ in $E^{n}$ is extended to $\hat{\sigma}:\hat{E}^{n}\rightarrow\hat{E}^{n}$ by setting $\hat{\sigma}(\infty)=\infty.$ The extended map $\hat{\rho}$ (resp. $\hat{\sigma}$) is called the reflection of $\hat{P}(a,t)$ (resp. $\hat{S}(a,r)$) , which is a homeomorphism of $\hat{E}^{n}$.
Definition 3.7.
A sphere of $\hat{E}^{n}$ is either an extended hyperplane or an Euclidean sphere. A Möbius transformation of $\hat{E}^{n}$ is a finite composition of reflection of a sphere in $\hat{E}^{n}$.
3.3. The conformal ball model
In this section, we recall the conformal ball model of hyperbolic space. Identify $\mathbb{R}^{n}$ with $\mathbb{R}^{n}\times\{0\}$ in $\mathbb{R}^{n+1}$. Let $B^{n}$ be the open unit ball of $\mathbb{R}^{n},$ and let $|x|$ be the Euclidean norm of a vector $x$ in $\mathbb{R}^{n}$: i.e., $|x|=(x_{1}^{2}+\cdots+x_{n}^{2})^{\frac{1}{2}}.$
The stereographic projection $\zeta$ from $B^{n}$ to the hyperbolic space $H^{n}$ is defined by the formula
$\zeta(x)=\Bigl{(}\hskip 1.42262pt\frac{2x_{1}}{1-|x|^{2}},\cdots,\frac{2x_{n}}{1-|x|^{2}},\frac{1+|x|^{2}}{1-|x|^{2}}\hskip 1.42262pt\Bigl{)}.$
This map $\zeta$ is bijective.
Definition 3.8.
A metric $d_{B}$ on $B^{n}$ is defined by the formula $d_{B}(x,y)=d_{H}(\zeta(x),\zeta(y)).$ The metric space $(B^{n},d_{B})$ is called the conformal ball model of hyperbolic $n$-spaces.
Definition 3.9.
A Möbius transformation of $B^{n}$ is a Möbius transformation $\phi$ of $\hat{E}^{n}$ that leaves $B^{n}$ invariant: i.e. $\phi$ satisfies $\phi(B^{n})=B^{n}.$ The group of Möbius transformation of $B^{n}$ is denoted by $M(B^{n})$.
It is well-known that every Möbius transformation of $B^{n}$ restricts to an isometry of the conformal ball model $B^{n}$ and the restriction induces an isomorphism from $M(B^{n})$ to $I(B^{n}).$
The boundary of $B^{n}$ in $\mathbb{R}^{n}$, denoted by $\partial B^{n}$, is the unit sphere $\{x\in\mathbb{R}^{n}:|x|=1\}$, where $|x|$ is the Euclidean norm of $x$. A hyperbolic m-plane of $B^{n}$ is the image of a hyperbolic $m$-plane of $H^{n}$ by $\zeta^{-1}$. By [18, Theorem 4.5.3], a hyperbolic $m$-plane of $B^{n}$ is the intersection of $B^{n}$ with either an $m$-dimensional vector subspace of $\mathbb{R}^{n}$ or an $m$-dimensional sphere of $\mathbb{R}^{n}$ orthogonal to $\partial B^{n}.$
Definition 3.10.
Let $c$ be a point in $\partial B^{n}$. Let $S$ be an $(n-1)$-dimensional inscribed sphere of $B^{n}$ tangent at $c$, and let $T$ be the one of connected component of $\mathbb{R}^{n}\backslash S$ which is compact. We call the intersection of $B^{n}$ with $S$ (resp. $T$) the horosphere of $B^{n}$ (resp. horoball of $B^{n}$).
3.4. The upper half-space model
Definition 3.11.
Let $x=(x_{1},\cdots,x_{n})$ be a vector in $\mathbb{R}^{n}$, and let $e_{n}$ be the vector in $\mathbb{R}^{n}$ whose entries are all zero except for the $n$-th, which is one. Let $\rho$ (resp. $\sigma$) be the reflection of $\hat{E}^{n}$ in $P(e_{n},0)$ (resp. $S(e_{n},\sqrt{2})$). The maps $\rho$ and $\sigma$ is given by the formula
$\rho(x_{1},\cdots,x_{n})=(x_{1},\cdots,-x_{n}),$
$\sigma(x)=e_{n}+\frac{2(x-e_{n})}{|x-e_{n}|^{2}}{\text{ for $x\in\mathbb{R}^{n}\backslash\{e_{n}\}$}}.$
The $n$-dimensional upper half-space $U^{n}$ is defined to be
$U^{n}=\{(x_{1},\cdots,x_{n})\in\mathbb{R}^{n}:x_{n}>0\}.$
Then the map $\eta=\sigma\rho$ is a map from the upper half-space $U^{n}$ to the open unit ball $B^{n}$, called the standard transformation from $U^{n}$ to $B^{n}.$ The Poincaré metric on $U^{n}$, denoted by $d_{U}$ is defined by the formula
$d_{U}(x,y)=d_{B}(\eta(x),\eta(y)).$
The pair $(U^{n},d_{U})$ is a metric space, called the upper half-model of hyperbolic $n$-spaces. By the definition, this is isomorphic to the conformal ball model $B^{n}$ of hyperbolic $n$-spaces. The group of isometries of $U^{n}$ is denoted by $I(U^{n})$.
Definition 3.12.
A Möbius transformation of $U^{n}$ is a Möbius transformation $\phi$ that leaves $U^{n}$ invariant: i.e. $\phi$ satisfies $\phi(U^{n})=U^{n}.$ The group of Möbius transformation of $U^{n}$ is denoted by $M(U^{n}).$
Every Möbius transformation of the upper half-space model $U^{n}$ restricts to an isometry of $U^{n}$ and the restriction induces an isomorphism from $M(U^{n})$ to $I(U^{n}).$
Definition 3.13.
Identify $E^{n-1}$ with $E^{n-1}\times\{0\}$ in $E^{n}.$ Let $\phi$ be a Möbius transformation of $\hat{E}^{n-1}$, and let $\overline{a}=(a,0)$. The extension $\overline{\phi}$ of $\phi$ is defined as follows:
  1. (1)
    If $\phi$ is the reflection of $\hat{E}^{n-1}$ in $P(a,t)$, then $\overline{\phi}$ is the reflection of $\hat{E}^{n}$ in $P(\overline{a},t)$.
  2. (2)
    If $\phi$ is the reflection of $\hat{E}^{n-1}$ in $S(a,t)$, then $\overline{\phi}$ is the reflection of $\hat{E}^{n}$ in $S(\overline{a},t)$.
  3. (3)
    In the general case, $\phi$ is the composition $\phi=\sigma_{1}\cdots\sigma_{m}$ of reflections. Then, we define the map $\overline{\phi}$ by the composition $\overline{\phi}=\overline{\sigma_{1}}\cdots\overline{\sigma_{m}}.$
We call the extension of $\phi$ the Poincaré extension
Poincaré extension is well-defined. Moreover, the following holds:
Theorem 3.14.
A Möbius transformation $\phi$ of $\hat{E}^{n}$ leaves $U^{n}$ invariant if and only if $\phi$ is the Poincaré extension of a Möbius transformation of $\hat{E}^{n-1}.$ In particular, Poincaré extension induces the isomorphism from $M(\hat{E}^{n-1})$ to $M(U^{n})$.
Definition 3.15.
Let $\eta$ be the standard transformation from $U^{n}$ to $B^{n}$. A subset $P$ is a hyperbolic $m$-plane (resp. horosphere, horoball) of $U^{n}$ if the image $\eta(P)$ of $P$ by $\eta$ is a hyperbolic $m$-plane (resp. horosphere, horoball) in $B^{n}$.
Theorem 3.16 ( [18, Theorem 4.6.3. and 4.6.5.] ).
  1. (1)
    A subset $P$ of $U^{n}$ is a hyperbolic $m$-plane of $U^{n}$ if and only if $P$ is the intersection of $U^{n}$ with either an $m$-plane of $E^{n}$ orthogonal to $E^{n-1}$ or an $m$-sphere of $E^{n}$ orthogonal to $E^{n-1}$.
  2. (2)
    A subset $\Sigma$ of $U^{n}$ is a horopshere of $U^{n}$ based at a point $b$ of $\hat{E}^{n-1}$ if and only if $\Sigma$ is either a Euclidean hyperplane in $U^{n}$ parallel to $E^{n-1}$ if $b=\infty,$ or the intersection with $U^{n}$ of a Euclidean sphere in $\overline{U}^{n}$ tangent to $E^{n-1}$ at $b$ if $b\neq\infty.$
3.5. Types of transformations and elementary groups
Let $B^{n}$ be the conformal ball model of $n$-dimensional hyperbolic spaces. We write $\overline{B^{n}}$ for the closed unit ball $B^{n}\cup\partial B^{n}.$
Definition 3.17.
Let $\phi$ be an element of $M(B^{n})$. Then $\phi$ maps the closed unit ball $\overline{B^{n}}$ to itself. By the Brouwer fixed point theorem, $\phi$ has a fixed point in $\overline{B^{n}}$.
  1. (1)
    $\phi$ is elliptic if $\phi$ has a fixed point in $B^{n}$,
  2. (2)
    $\phi$ is parabolic if $\phi$ has no fixed point in $B^{n}$ and has a unique fixed point in $\partial B^{n}$,
  3. (3)
    $\phi$ is hyperbolic if $\phi$ has no fixed point in $B^{n}$ and fixes two points in $\partial B^{n}$.
Let $\phi$ be an isometry of $U^{n}$, and let $\eta$ be the standard transformation from $U^{n}$ to $B^{n}$. Then $\phi$ is said to be elliptic (resp. parabolic, hyperbolic) if $\eta\phi\eta^{-1}$ is elliptic (resp. parabolic, hyperbolic).
Definition 3.18.
A subgroup $G$ of $M(B^{n})$ is elementary if $G$ has a finite orbit in the closed ball $B^{n}$. Elementary groups can also be classified into three types as follows:
  1. (1)
    The group $G$ is of elliptic type if $G$ has a finite orbit in $B^{n}$.
  2. (2)
    The group $G$ is of parabolic type if $G$ fixes a point of $\partial B^{n}$ and has no other finite orbit in $\overline{B}^{n}$
  3. (3)
    The group $G$ is of hyperbolic type if $G$ is neither of elliptic type nor parabolic type.
Theorem 3.19 ( [18, Theorem 5.5.2.] ).
Let $\Gamma$ be a subgroup of $M(B^{n})$. Then $\Gamma$ is an elementary discrete subgroup of elliptic type if and only if $\Gamma$ is conjugate in $M(B^{n}$ to a finite subgroup of the orthogonal group $O(n)$.
Theorem 3.20 ( [18, Theorem 5.5.5] ).
Let $\Gamma$ be a subgroup of $M(U^{n}).$ The group $\Gamma$ is an elementary discrete subgroup of parabolic type if and only if $\Gamma$ conjugate in $M(U^{n})$ to an infinite discrete subgroup of $I(E^{n-1}).$
3.6. Dirichlet domains
Definition 3.21.
Let $X$ be a metric space. A subset $D$ of $X$ is a fundamental domain for a group $\Gamma$ of isometries of $X$ if
  1. (1)
    the set $D$ is open and connected in $X$,
  2. (2)
    the elements of $\{gD:g\in\Gamma\}$ are mutually disjoint,
  3. (3)
    $X=\bigcup_{g\in\Gamma}\hskip 2.84526ptg\overline{D}$,
    where $\overline{D}$ is the closure of $D$ in $X$.
A fundamental domain $D$ is locally finite if and only if the set $\{g\overline{D}:g\in\Gamma\}$ is locally finite: i.e., for each $x\in X$, there is an $\epsilon>0$ such that the open ball $B(x,\epsilon)$ of radius $\epsilon$ centered at $x$ meets only finitely many members of $\{g\overline{D}:g\in\Gamma\}$.
Let $\Gamma$ be a discontinuous group of isometries of a hyperbolic $n$-space $H^{n}$. By [18, Theorem 6.6.10 and Theorem 6.6.12], there is a point $a$ such that the stabilizer group $\Gamma_{a}$ of $a$ is trivial. For each $g\neq 1$ in $\Gamma$, we set $H_{g}^{+}(a)=\{x\in H^{n}:d_{H}(x,a)<d_{H}(x,ga)\}.$
Definition 3.22.
The Dirichlet domain $D_{\Gamma}(a)$ for $\Gamma$, with center $a$, is either $H^{n}$ if $\Gamma$ is trivial or
$D_{\Gamma}(a)=\bigcap_{g\in\Gamma\backslash\{1\}}H_{g}^{+}(a).$
Let $D_{\Gamma}(a)$ be the Dirichlet domain for $\Gamma$ with center $a$ as above. By [18, Theorem 6.6.13], $D_{\Gamma}(a)$ is a locally finite fundamental domain for $\Gamma$.
3.7. Limit sets
Definition 3.23.
A point $a$ of $\partial B^{n}$ is a limit point of a subgroup $G$ of $M(B^{n})$ if there is a point $x\in B^{n}$ and a sequence $\{g_{i}\}$ of elements of $G$ such that the sequence $\{g_{i}x\}_{i\in\mathbb{N}}$ converges to $a$. The set of limit points of $G$ is called the limit set of $G$, denoted by $\Lambda(G)$. The complement of $\Lambda(G)$ in $\partial B^{n}$ is called the ordinary set of $G$, denoted by $O(G).$
Definition 3.24.
A point $a$ of $\partial B^{n}$ is a conical limit point of a subgroup $G$ of $M(B^{n})$ if there is a point $x\in B^{n}$, a sequence $\{g_{i}\}$ of elements of $G$, a $1$-plane $R$ and a real number $r>0$ such that the closure of $R$ in $\overline{B^{n}}$ contains $a$, $g_{i}x$ is in the $r$-neightborhood $N(R,r)$ of $R$ and the sequence $\{g_{i}x\}_{i\in\mathbb{N}}$ converges to $a$, where $N(R,r)$ is defined to be
$N(R,r)=\{x\in B^{n}:d_{B}(x,y)<r{\text{ for some }}y\in R\}.$
Definition 3.25.
A point $a$ of $\partial B^{n}$ is a bounded parabolic limit point of a discrete group $\Gamma$ of $M(B^{n})$ if $a$ is a fixed point of a parabolic element of $\Gamma$ and the orbit space $(\Lambda(G)\backslash\{a\})/\Gamma$ is compact.
3.8. Geometrically finite discrete groups
Definition 3.26.
A convex polyhedron $P$ in $B^{n}$ is geometrically finite if for any $x\in\overline{P}\cap\partial B^{n-1},$ there exists an open neighborhood $N$ of $x$ in $E^{n}$ such that $N$ meets only the sides of $P$ whose closure in $E^{n}$ contains $x$.
Definition 3.27.
A discrete group $\Gamma$ of $M(B^{n})$ is geometrically finite if $\Gamma$ has a geometrically finite convex fundamental polyhedron $P$ and $P$ is exact: i.e., for each side $S$ of $P$, there is an element $\gamma$ of $\Gamma$ such that $S=P\cap\gamma P.$
Definition 3.28.
Let $\Gamma$ be a discrete subgroups of $M(B^{n})$ with a parabolic element that fixes a point $a$ of $\partial B^{n}$. Let $\Gamma_{a}$ be the stabilizer group of $a$. A horoball $B(a)$ based at $a$ is said to be a horocusped region for $\Gamma$ based at $a$ if and only if for each $g\in\Gamma\backslash\Gamma_{a}$, we have
$B(a)\cap gB(a)=\emptyset.$
If a horocusped region $B(a)$ based at $a$ is not maximal with respect to the inclusion relation, $B(a)$ is said to be proper.
Definition 3.29.
Let $C(\Gamma)$ be the hyperbolic convex hull of the limit set $\Lambda(\Gamma)$ of $\Gamma$. The convex core of $M=B^{n}/\Gamma$ is the set
$C(M)=(C(\Gamma)\cap B^{n})/\Gamma.$
Theorem 3.30 ([18, Theorem 12.4.5.]).
Let $\Gamma$ be a discrete subgroup of $M(B^{n})$, let $M=B^{n}/\Gamma$, and let $C(M)$ be the convex core of $M$. The following are equivalent:
  1. (1)
    The group $\Gamma$ is geometrically finite,
  2. (2)
    There is a (possibly empty) finite union $V$ of proper horocusps of $M$, with disjoint closures, such that $C(M)\backslash V$ is compact, or
  3. (3)
    Every limit point of $\Gamma$ is either conical or bounded parabolic,
  4. (4)
    Every ecaxt convex fundamental polyhedron for $\Gamma$ is geometrically finite.
Theorem 3.31 ( [18, Theorem 12.4.10] ).
Let $\Gamma$ be a discrete subgroup of $M(B^{n})$ with $n>1$. Then the following are equivalent:
  1. (1)
    The orbit space $B^{n}/\Gamma$ has finite volume.
  2. (2)
    The group $\Gamma$ is geometrically finite and $\Lambda(\Gamma)=\partial B^{n}$.
  3. (3)
    There are finite points $p_{1},\cdots p_{k}$ of $\overline{B^{n}}$ such that $P\cap B^{n}$ is a fundamental domain with finite volume, where $P$ is the hyperbolic convex full of $\{p_{1},\cdots,p_{k}\}$ in $\overline{B^{n}}.$
Theorem 3.32 ( [18, Theorem 12.6.5] ).
Let $\Gamma$ be a subgroup of $M(B^{n})$, and let $C$ be the set of cusped limit points of $\Gamma$. For each $c\in C$, there exists a cusped region $U(c)$ based at $c$ for $\Gamma$ such that
  1. (1)
    $\{\overline{U}(c):c\in C\}$ is mutually disjoint,
  2. (2)
    $\gamma U(c)=U(\gamma c)$ for any $\gamma\in\Gamma$ and $c\in C$,
  3. (3)
    $\{\overline{U}(c)\backslash\{c\}\}$ is a locally finite family of closed subsets of $B^{n}\cup O(\Gamma).$
4. Cohomological dimension of groups
4.1. Cohomology of groups
In this section, we review some definitions and basic facts about cohomology of groups, which are contained in [8].
Definition 4.1.
Let $R$ be a ring and $M$ be a left $R$-module. Let $n$ is a non-negative integer. Then $\operatorname{prodim}_{R}M\leq n$ if and only if there is a left projective resolution of length $n$:
$0\rightarrow P_{n}\rightarrow P_{n-1}\rightarrow\cdots\rightarrow P_{0}\rightarrow M\rightarrow 0$
The following lemma is well-known.
Lemma 4.2 ([8] Chapter VIII, Lemma (2.1)).
The following conditions are equivalent:
  1. (1)
    $\operatorname{projdim}_{R}M\leq n.$
  2. (2)
    $\operatorname{Ext}_{R}^{i}(M,-)=0$ for $i>n$.
  3. (3)
    $\operatorname{Ext}_{R}^{n+1}(M,-)=0.$
  4. (4)
    For any left $R$-module exact sequence $0\rightarrow K\rightarrow P_{n-1}\rightarrow\cdots\rightarrow P_{0}\rightarrow 0$ with $P_{i}$ projective, $K$ is projective.
Definition 4.3.
We define the cohomological dimension $\operatorname{cd}\Gamma$ as following:
$\begin{split}\operatorname{cd}\Gamma&=\operatorname{projdim}_{\mathbb{Z}\Gamma}\mathbb{Z}\\ &=\inf\{n:\mathbb{Z}{\textup{ admits a projective resolution of lengh }}n\}\\ &=\inf\{n:H^{n}(\Gamma,-)=0{\textup{ for }}\forall i>n\}\\ &=\sup\{n:H^{n}(\Gamma,-)\neq 0{\textup{ for some }}\Gamma{\textup{-module }}M\}\end{split}$
Recall that the functor $\operatorname{Ext}_{R}^{i}(M,-)=H^{i}(\operatorname{Hom}_{R}(M,-))$. Then $\operatorname{Ext}_{\mathbb{Z}\Gamma}^{i}(M,-)=H^{i}(\Gamma,-)$. By the previous Lemma, the above definition is well-defined.
The following theorem is called Serre’s theorem that claims well-definedness of the virtual cohomological dimension.
Theorem 4.4 (Serre’s theorem, [8] Chapter VIII, Theorem (3.1)).
If $\Gamma$ is a torsion-free group and $\Gamma^{{}^{\prime}}$ is a subgroup of finite index, then $\operatorname{cd}\Gamma=\operatorname{cd}\Gamma^{{}^{\prime}}$.
Definition 4.5.
Let $\Gamma$ be a virtually torsion-free group. By the Serre’s theorem, any torsion-free subgroup of finite index has the same cohomological dimension. Then the dimension is called the virtual cohomological dimension, denoted by $\operatorname{vcd}\Gamma$.
4.2. Bestvina-Mess formula
In [5], Bestvina and Mess proved the following formula for any virtually torsion free hyperbolic group $G$:
$\operatorname{vcd}G=\dim\partial G+1,$
where $\partial G$ is the Gromov boundary of $G$ and $\dim\partial G$ is the covering dimension of $\partial G$. In [5], this formula, called the Bestvina-Mess formula, is generalized for groups which admits a $\mathcal{Z}$-structure. Recently, Fukaya [10] proved the Bestvina-mess formula for relatively hyperbolic groups.
Definition 4.6.
Let $A$ be a subset of a topological space $X$. If there exists a map $r:X\rightarrow A$ such that $r|A=id$, a subset A is called a retract of $X$. A neighborhood retract of $X$ is a closed set in $X$ that is a retract of some neighborhood in $X$.
Definition 4.7.
A metrizable space $X$ is called an abusolute neighborhood retract (ANR) (resp. an abusolute retract (AR)) if $X$ is a neighborhood retract (resp. a retract) of an arbitrary metrizable space that contains $X$ as a closed subspace.
It is well-known that a ANR space $X$ is AR if and only if $X$ is contractible.
Example 4.8.
1. Every topological manifold is ANR.
2. $[0,1]^{n}$ is AR.
Definition 4.9.
Let $X$ be a ANR space. A closed subset $Z$ of $X$ is called a $\mathcal{Z}$-set if $X\backslash Z$ is homotopy dense in $X$, i. e., there exists a homotopy $H:X\times[0,1]:\rightarrow X$ such that for any $t>0$, $H_{t}(X)\subset A$.
Definition 4.10.
Let $\overline{X}$ be a compact AR space, and let $Z$ be a closed subset of $\overline{X}$. A pair $(\overline{X},Z)$ is a $\mathcal{Z}$-structure for a group $G$ if
  1. (1)
    $Z$ is a $\mathcal{Z}$-set of $\overline{X}$.
  2. (2)
    There exists a proper metric $d$ on $X=\overline{X}\backslash Z$ such that 2-1. $G$ geometrically acts on $(X,d).$ 2-2. For any compact subset $K$ of $X$ and any open covering $\mathcal{U}$ of $\overline{X}$, there exists $U_{g}\in\mathcal{U}$ such that $gK\subset U_{g}$ for all $g\in G$ exept finitely many elements of $G$.
Theorem 4.11 ( [6, Theorem 1.7]).
Let $(\overline{X},Z)$ be a $\mathcal{Z}$-structure on a group $G$. Then $\operatorname{vcd}G=\dim Z+1,$ where $\dim Z$ is the covering dimension of $Z$.
Example 4.12.
Let $\mathbb{B}^{n}$ be the $n$-dimensional conformal model of hyperbolic spaces. and let $\Gamma$ be a cocompact discrete subgroup of $M(B^{n}).$ Then the pair (¯$(\overline{\mathbb{H}}^{n},S^{n-1})$ is a $\mathcal{Z}$-structure of a group $\Gamma$, and so $\operatorname{vcd}\Gamma=\dim S^{n-1}+1=n.$
5. Automorphism groups of K3 surfaces and its virtual cohomological dimension
In this section, we consider the natural actions of the automorphism groups of K3 surfaces on hyperbolic spaces. First, we recall the constructions of hyperbolic spaces induced by hyperbolic lattices, and discuss the actions of the automorphism groups of K3 surfaces on certain hyperbolic spaces, and moreover some properties of automorphism groups. Secondly, we discuss some invariants of automorphism groups, for example the virtual cohomological dimension.
5.1. Geometrical finiteness of $\operatorname{Aut}_{s}(X)$
Let $X$ be an algebraic K3 surface defined over $\mathbb{C}$, and let $f$ be the quadratic form defined by $f(x)=\langle x,x\rangle$ for $x\in NS(X)\otimes\mathbb{R}$, where $\langle$,$\rangle$ , ⟩ is the inner product on $NS(X)$ induced by the cup product on $H^{2}(X,\mathbb{Z})$. $f$ has the signature $(1,n)$, where $n=\rho(X)-1.$ Then there exists $M\in\mathrm{GL}(n+1,\mathbb{R})$ such that $f_{n+1}(Mx)=f(x)$, where $f_{n+1}(x)=x_{0}^{2}-(x_{1}^{2}+x_{2}^{2}+\cdots+x_{n}^{2}).$ Therefore, we can identify the component of the set
$\{x\in NS(X)\otimes\mathbb{R}:\langle x,x\rangle=1\}$
containing ample classes of $X$ with the hyperboloid model of hyperbolic $n$-space. We denote this hyperboloid model of hyperbolic $n$-space by $\mathbb{H}^{n}_{X}$. The hyperbolic distance function $d$ on $\mathbb{H}^{n}_{X}$ satisfies $\cosh d(x,y)=\langle x,y\rangle.$
Definition 5.1.
We call $\mathbb{H}^{n}_{X}$ the hyperbolic space associated to $X$. Let $\overline{P^{+}}(X)$ be the closure of the ample cone $P^{+}(X)$ in $NS(X)\otimes\mathbb{R}.$ We define $\mathbb{P}(\overline{P^{+}}(X))$ by
$\mathbb{P}(\overline{P^{+}}(X))=(\overline{P^{+}}(X)\backslash\{0\})/\mathbb{R}_{>0}.$
For a point $x$ in $\overline{P^{+}}(X)\backslash\{0\}$, we denote by $[x]$ the class in $\mathbb{P}(\overline{P^{+}}(X))$ corresponding to $x$. The inclusion from $\mathbb{H}^{n}_{X}$ to $\overline{P^{+}}(X)$ induces the continuous injective map $i:\mathbb{H}^{n}_{X}\rightarrow\mathbb{P}(\overline{P^{+}}(X)).$ Since $i$ is a homeomorphism from $\mathbb{H}^{n}_{X}$ to $i(\mathbb{H}^{n}_{X})$, we identify $\mathbb{H}^{n}_{X}$ with $i(\mathbb{H}^{n}_{X}).$
The boundary of $\mathbb{H}^{n}_{X}$ at infinity, denoted by $\partial{\mathbb{H}}^{n}_{X}$, is the closure of $i(\mathbb{H}^{n}_{X})$ in $\mathbb{P}(\overline{P^{+}}(X)),$ which is equals to the set $\{x\in\overline{P^{+}}(X)\backslash\{0\}:\langle x,x\rangle=0\}/\mathbb{R}_{>0}.$ We set ¯$\overline{\mathbb{H}}^{n}_{X}=\mathbb{H}^{n}_{X}\cup\partial{\mathbb{H}}^{n}_{X}.$
Definition 5.2.
Let $A(X)$ be the ample cone. We define
$\operatorname{Aut}(A(X)):=\{\phi\in O(NS(X)):\phi A(X)=A(X)\}{\text{ and }}A_{X}:=A(X)\cap\mathbb{H}^{n}_{X}$
The closure of $A_{X}$ in $\overline{\mathbb{H}^{n}_{X}}$ is denoted by $\overline{A_{X}}.$ We denote $\overline{A_{X}}\backslash A_{X}$ by $\partial A_{X}$, called the boundary associated with the ample cone of $X$.
By the unimodulality of $H^{2}(X,\mathbb{Z})$, the discriminant groups of $NS(X)$ and $T_{X}$ are isomorphic:
$A_{NS(X)}=NS(X)^{*}/NS(X)\simeq T_{X}^{*}/T_{X}=A_{T_{X}}.$
Each isometry $\phi$ of $NS(X)$ or $T_{X}$ gives rise to an isometry of the group $A_{NS(X)}\simeq A_{T(X)}$, denoted by $\phi^{*}$. Since the automorphism group $\operatorname{Aut}(X)$ of $X$ naturally acts on the ample cone of $X$, there is a natural homomorphism $\rho:\operatorname{Aut}(X)\rightarrow\operatorname{Aut}(A(X)).$ As for this homomorphism, the following theorem is well-known.
Theorem 5.3 ([13] Theorem 8.1).
The kernel and the cokernel of $\rho$ are finite groups.
Theorem 5.4 ([21] Proposition 2.2).
Let $X$ be an algebraic K3 surface over $\mathbb{C}$. Then $\operatorname{Aut}(X)$ is finitely generated.
An elementary algebraic proof of the following lemma can be found in [1].
Lemma 5.5 (Selberg’s lemma, [1]).
Let $K$ be a field of characteristic zero. Any finitely generated subgroup of $\operatorname{GL}(n,K)$ is virtually torsion-free. In other words, such a group has a torsion free subgroup of finite index.
Definition 5.6.
Let $X$ be a K3 surface, and let $\omega_{X}$ be a nowhere vanishing holomorphic 2-form of $X$. A automorphism $\phi$ of $X$ is said to be symplectic if $\phi$ acts trivially on $\omega_{x}$. The symplectic automorphism group of $X$, denoted by $\operatorname{Aut}_{s}(X)$, is the group of symplectic automorphisms of $X$.
The following lemma is well-known.
Lemma 5.7 ([13, Lemma 8.11]).
The symplectic automorphism of $X$ acts trivially on the transcendental lattice $T_{X}$.
Let $\phi$ be the symplectic automorphism of $X$. Then, by Lemma 5.7, $\phi$ acts trivially on $A_{NS(X)}\simeq A_{T_{X}}.$ By the Torelli theorem, we can identify $\operatorname{Aut}_{s}(X)$ with the subgroup $\{\phi\in O^{+}(NS(X)):\phi A(X)=A(X),\phi^{*}|_{A_{NS(X)}}=\mathrm{id}_{A_{NS(X)}}\}$. Therefore, $\operatorname{Aut}_{s}(X)$ is a subgroup of finite index in $\operatorname{Aut}(A(X)).$ Since any element of $O^{+}(NS(X))$ is an isometry of the hyperbolic space $\mathbb{H}^{n}_{X}$ associated $X$, the symplectic automorphism group $\operatorname{Aut}_{s}(X)$ of $X$ is an $n$-dimensional discrete subgroup of isometries of $\mathbb{H}^{n}_{X}$.
Proposition 5.8.
Let $X$ be a K3 surface, and let $a\in A_{X}.$ Then the Dirichlet domain $D(a)$ of $\mathbb{H}^{n}_{X}$, with center $a$, for the reflection group $W(X)$ is equal to $A_{X}=A(X)\cap\mathbb{H}^{n}_{X}.$
Proof.
Recall that $H^{+}_{w}(a)=\{x\in\mathbb{H}^{n}:d(x,a)<d(x,wa)\hskip 2.84526pt\}$ for $w\in W_{X}.$ If $w$ is the reflection $s_{\delta}$ for some $(-2)$-curve $\delta$,
$\hskip 19.91692ptd(x,a)<d(x,s_{\delta}a)$
$\Leftrightarrow\hskip 2.84526pt\cosh d(x,a)<\cosh d(x,s_{\delta}a)$
$\Leftrightarrow\hskip 2.84526pt\langle x,a\rangle<\langle x,s_{\delta}a\rangle$
$\Leftrightarrow\hskip 5.69054pt0<\langle a,\delta\rangle\langle x,\delta\rangle$
$\Leftrightarrow\hskip 5.69054pt0<\langle x,\delta\rangle$
Then we have that $H^{+}_{s_{\delta}}(a)=\{x\in\mathbb{H}^{n}_{X}:\langle x,\delta\rangle>0\}.$ Since $A(X)$ is a fundamental domain for $W(X)$ on the positive cone $P^{+}(X)$ of $X$ whose sides are bounded by $(-2)$-curves, we have $A_{X}=\cap_{\delta\in\Delta^{+}(X)}H^{+}_{s_{\delta}}.$
By the definition of $D(a)$, we have $D(a)\subset A_{X}.$ Since $D(a)$ and $A_{X}$ are fundamental domains for $W(X)$, we have $D(a)=A_{X}.$
$O^{+}(NS(X))$ is a geometrically finite subgroup of the first kind. By Theorem 3.31, we have the following:
Lemma 5.9.
Let $X$ be a K3 surface. Then a fundamental domain of $\mathbb{H}^{n}_{X}$ for $O^{+}(NS(X))$ is a generalized polytope. In particular, the action of $O^{+}(NS(X))$ on $\mathbb{H}^{n}_{X}$ is finite covolume.
Lemma 5.10.
Let $a\in A_{X}$, and let $P$ be the Dirichlet domain, centered at $a$, of $\mathbb{H}^{n}_{X}$ for $O^{+}(NS(X))$. Then, $\operatorname{Aut}(A(X))\cdot\overline{P}=\overline{A_{X}}$, where $\overline{P}$ (resp. $\overline{A_{X}}^{H}$) are the closure of $P$ (resp. $A_{X}$) in $\mathbb{H}^{n}_{X}.$
Proof.
By Lemma 5.8 and the definition of $P$, we have
$P=\{x\in\mathbb{H}^{n}_{X}:d(x,a)<d(x,ga){\text{ for any }}g\in O^{+}(NS(X))\},$
$A_{X}=\{x\in\mathbb{H}^{n}_{X}:d(x,a)<d(x,wa){\text{ for any }}w\in W(X)\}.$
Then we have $P\subset A_{X}$ and so $\overline{P}\subset\overline{A_{X}}^{H}.$ Hence we obtain $\operatorname{Aut}(A(X))\cdot\overline{P}\subset\overline{A_{X}}^{H}.$
Conversely, because the tessellation $\{g\overline{P}:g\in O^{+}(NS(X))\}$ of $\mathbb{H}^{n}_{X}$ is locally finite, the union $\operatorname{Aut}(A(X))\cdot\overline{P}$ is a closed subset. Then it is enough to show that $A_{X}\subset\operatorname{Aut}(A(X))\cdot\overline{P}.$ For any $x\in A_{X}$, there exist $w\in W(X)$ and $\gamma\in\operatorname{Aut}(A(X))$ such that $x\in w\gamma\overline{P}.$ Then we have $x\in w\gamma\overline{P}\cap A_{X}\subset w\overline{A_{X}}^{H}\cap A_{X}.$ By [18] Theorem 6.6.4, we have $w=1$. In particular, we have $x\in\gamma\overline{P}$. Hence we have $A_{X}\subset\operatorname{Aut}(A(X))\cdot\overline{P}$. ∎
Definition 5.11.
Let $P$ be the same as in Lemma 5.10. We define sets of bounded parabolic points of $O^{+}(NS(X))$ as follows:
  • $I_{P}=$ the set of ideal vertices of $P$,
  • $V=Aut(A(X))\cdot I_{P}$,
  • $V_{NS}=O^{+}(NS(X))\cdot I_{P},$
  • $C_{X}=\{c\in V:\operatorname{vcd}W(X)_{c}<n-1\}.$
Definition 5.12.
Let $X$ be a K3 surface. Since $\operatorname{Aut}_{s}(X)$ is of finite index in $\operatorname{Aut}(A(X))$, the limit set of $\operatorname{Aut}_{s}(X)$ is equal to that of $\operatorname{Aut}(A(X))$. We denote this limit set by $\Lambda_{X}$, called the limit set of the K3 surface $X$. We denote the hyperbolic convex hull of $\Lambda_{X}$ by $C(X)$, and put $B(X)=C(X)\cap\mathbb{H}^{n}_{X}$.
Lemma 5.13.
For any $c\in V\backslash C_{X}$, there exists a horoball $U(c)$ based at $c$ such that $B(X)\cap U(c)=\varnothing$. Moreover, the point $c$ is isolated in the boundary $\partial A_{X}$ associated with the ample cone.
Proof.
If $\operatorname{Aut}(A(X))$ is elementary, this lemma is obvious.
Suppose that $\operatorname{Aut}(A(X))$ is non-elementary. Let $\phi:\mathbb{H}^{n}_{X}\rightarrow U^{n}$ be an isometry from $\mathbb{H}^{n}_{X}$ to the upper half-space model $U^{n}$ such that $\phi(c)=\infty.$ The isometry $\phi$ induces the natural homeomorphisms from $\partial\mathbb{H}^{n}_{X}$ to $\partial U^{n}.$ We also denote this homeomorphism by $\phi$. Let $b$ be a point of $U^{n}$ corresponding to a point $a$ in $A_{X}$. We can consider the stabilizer group $W(X)_{\infty}$ as the Poincaré extension of a subgroup of $I(E^{n-1})$. Since $\operatorname{vcd}W(X)_{c}=n-1$, $W(X)_{\infty}$ is a crystallographic group of $E^{n-1}$. Then the Dirichlet domain $K$ of $E^{n-1}$ for $W(X)_{\infty}$, centered at $b$, is compact. Proposition 5.8 implies that $\Lambda_{X}\subset\partial A_{X}.$ Then we have $\phi(\Lambda_{X})\subset K\cup\{\infty\}.$ Since $\operatorname{Aut}(A(X))$ is non-elementary, the limit set $\Lambda_{X}$ is perfect: i.e., every point of $\Lambda_{X}$ is not isolated. Hence the infinity $\infty$ is not in $\phi(\Lambda_{X})$. By the compactaness of $K$, there exists a horoball $B_{\infty}$ based at $\infty$ such that $C(K)\cap B_{\infty}=\varnothing,$ where $C(K)$ is the hyperbolic convex hull of $K$. In particular, $B(X)\cap B_{\infty}=\varnothing.$ The inverse image $\phi^{-1}(B_{\infty})$ of $B_{\infty}$ by $\phi$ is the desired horoball. ∎
Lemma 5.14.
Let $e$ be an elliptic divisor of $X$, let $f_{e}$ be an elliptic fibration associated to $e$, and let $\operatorname{Aut}(X)_{e}$ be the stabilizer group of $e$. Then
$\operatorname{vcd{\operatorname{Aut}(X)_{e}}}=\operatorname{vcd}\operatorname{MW}(f_{e}){\text{, and }}\operatorname{vcd}W(X)_{e}=\sum_{v\in\mathbb{P}^{1}}{(m_{v}-1)},$
where $\mathrm{MW}(f_{e})$ is the Mordell-Weil group of $f_{e}$, and $m_{v}$ be the number of connected components of the fiber $f^{-1}(v)$ for $v\in\mathbb{P}^{1}.$
Proof.
Let $v$ be a point in $\mathbb{P}^{1},$ and let $m_{v}$ be the number of connected components of the fiber $f^{-1}(v)$. By Shioda-Tate formula,
$\rho(X)-2=\operatorname{vcd}{MW(f_{e})}+\sum_{v\in\mathbb{P}^{1}}{(m_{v}-1)}.$
If $m_{v}\neq 1$, we denote by $W_{v}$ the Weyl group generated by the connected components of $f^{-1}(v)$, which is of type $\tilde{A}$, $\tilde{D}$, or $\tilde{E}.$ Since $W_{v}$ fixes the elliptic divisor $e$, we have $\oplus W_{v}\subset W(X)_{e}.$ Here, for each $v$ with $m_{v}\neq 1$, the Weyl group $W_{v}$ acts on the $(m_{v}-1)$-dimensional Euclidean space, and admits an $(m_{v}-1)$-dimensional simplex $\Delta_{v}$ as a fundamental domain for the action of $W_{v}.$ Hence $\oplus W_{v}$ acts on the $(\sum_{v\in\mathbb{P}^{1}}{(m_{v}-1})\big{)}$-dimensional Euclidean space, and has the fundamental domain $\prod_{v}{\Delta_{v}}$. Then we have
$\sum_{v\in\mathbb{P}^{1}}(m_{v}-1)=\operatorname{vcd}(\oplus W_{v})\leq\operatorname{vcd}W(X)_{e}.$
Since the Mordell-Weil group of $f_{e}$ is a subgroup of $\operatorname{Aut}(X)_{e}$, then $\operatorname{rk}\operatorname{MW}f_{e}\leq\operatorname{vcd}\operatorname{Aut}(X)_{e}$, where $\operatorname{rk}\operatorname{MW}f_{e}$ is the rank of Mordell-Weil group of $f_{e}$.
Since both $W(X)_{e}$ and $\operatorname{Aut}(X)_{e}$ fix the elliptic divisor $e$, which is a cusped limit points of $O^{+}(NS(X))$, these two groups are elementary of either elliptic or parabolic type. Therefore, there exists an integer $l$ (resp. $m$) such that $W(X)_{e}$ (resp. $\operatorname{Aut}(X)_{e}$) has a subgroup that is isomoprhic to $\mathbb{Z}^{l}$ (resp. $\mathbb{Z}^{m}$) of finite index in $W(X)_{e}$ (resp. $\operatorname{Aut}(X)_{e}$). Hence we have
(5.1) $\sum_{v\in\mathbb{P}^{1}}{(m_{v}-1)}\leq l,$
(5.2) $\operatorname{rk}\operatorname{MW}f_{e}\leq m,$
so that $\rho(X)-2\leq l+m.$ On the other hand, $O^{+}(NS(X))_{e}$ has a subgroup that is isomorphic to $\mathbb{Z}^{l+m}.$ Then $l+m=\rho(X)-2.$ The inequality 6.2.1 and 6.2.2 imply that
$l=\sum_{v\in\mathbb{P}^{1}}(m_{v}-1),$
$m=\operatorname{rk}\operatorname{MW}f_{e}.$
These are namely the desired equalities. ∎
By Lemma 5.14, we have $C_{X}=\{c\in V:\operatorname{vcd}\operatorname{Aut}(X)_{c}>1\}=\{c\in V:\operatorname{vcd}\operatorname{Aut}(A(X))_{c}>1\}.$ Then each element of $C_{X}$ is fixed by some parabolic transformation of $\operatorname{Aut}(A(X))_{c}.$
As noted in the introduction, the following theorem is also found by Kohei Kikuta at about the same time (see the introduction for details).
Theorem 5.15.
Let $X$ be a K3 surface, and let $\operatorname{Aut}_{s}(X)$ be the symplectic automorphism group of $X$. Then $\operatorname{Aut}_{s}(X)$ is geometrically finite.
Proof.
By Theorem 5.3, the symplectic automorphism group $\operatorname{Aut}_{s}(X)$ is a subgroup of finite index in $\operatorname{Aut}(A(X)).$ Then it is enough to show that $\operatorname{Aut}(A(X))$ is geometrically finite. Recall the notation of $P$, $V$, $V_{NS}$, and $C_{X}$ in Definition 5.11
By Theorem 3.32, there exist proper horoballs at $c\in V_{NS}$ such that
  1. (1)
    $\{\overline{U(c)}:c\in V_{NS}\}$ are mutually disjoint,
  2. (2)
    $gU(c)=U(gc)$ for any element $g\in O^{+}(NS(X))$,
  3. (3)
    $\{\overline{U(c)}\backslash{c}\}$ is locally finite in $\mathbb{H}^{n}_{X}$.
By Lemma 5.13, shrinking the horoballs $U(c)$ sufficiently small, we obtain that
$B(X)\backslash\bigcup_{c\in C_{X}}U(c)=B(X)\backslash\bigcup_{c\in V}U(c).$
By applying Lemma 5.10,
$B(X)\backslash\bigcup_{c\in C_{X}}U(c)$ $\subset\hskip 2.84526pt\overline{A_{X}}^{H}\backslash\bigcup_{c\in V}U(c)$
$=\bigl{(}\bigcup_{g\in\operatorname{Aut}(A(X))}g\overline{P}\hskip 2.84526pt\bigr{)}\big{\backslash}\bigcup_{c\in V}U(c)$
[…]
$=\bigcup_{g\in\operatorname{Aut}(A(X))}g(\overline{P}\backslash\bigcup_{c\in V}U(c)).$
Hence we have
(5.3) $B(X)\backslash\bigcup_{c\in C_{X}}U(c)\subset\bigcup_{g\in\operatorname{Aut}(A(X))}g(\overline{P}\backslash\bigcup_{c\in V}U(c)).$
Let $\pi:\mathbb{H}^{n}_{X}\rightarrow\mathbb{H}^{n}_{X}/\operatorname{Aut}(A(X))$ be the natural quotient map. We define $V_{c}=U(c)/\operatorname{Aut}(A(X)).$ By the equality 5.3, we obtain
$(B(X)/\operatorname{Aut}(A(X)))\backslash\bigcup_{c\in I_{P}\cap C_{X}}V_{c}\subset\pi(\overline{P}\backslash\bigcup_{c\in V}U(c)).$
Then $(B(X)/\operatorname{Aut}(A(X)))\backslash\bigcup_{c\in I_{P}\cap C_{X}}V_{c}$ is compact and $\{V_{c}\}_{c\in I_{P}\cap C_{X}}$ is a set of finitely many horocusps with disjoint closures. By Theorem 3.30, $\operatorname{Aut}(A(X))$ is geometrically finite. ∎
Proposition 5.16.
Let $p$ be a cusped limit point of $\operatorname{Aut}(A(X)).$ Then $p$ is a horopoint of $\overline{A_{X}},$ i.e., there exists a horoball $U$ based at $p$ such that $U$ meets just the sides of $\overline{A_{X}}$ incident with $p$.
Proof.
We now pass to the upper half-space model $U^{n}$ and conjugate $\operatorname{Aut}(A(X))$ so that $p=\infty.$ Let $P$ be a Dirichlet polyhedron $P$, centerd at $a\in A_{X}$, for $O^{+}(NS(X)).$ From the proof of [18, Theorem 12.4.4], there exist a horoball $\Sigma$ based at $\infty$ and elements $g_{1},\cdots,g_{k}$ in $O^{+}(NS(X))$ such that
$\Sigma\subset\bigcup_{f\in\operatorname{Aut}(A(X))_{\infty}}(\bigcup_{i=1}^{k}g_{i}P)$
and $\Sigma$ meets just the vertical sides of $fg_{i}P$.
Let $S$ be a side of $A_{X}$ with $S\cap\Sigma\neq\varnothing.$ There exists an element $x\in S$ and $r>0$ such that $B(x,r)\cap S$ is homeomorphic to $B^{n-1}.$ Since $P$ is locally finite, shrinking $r$, there exists $g_{1},\cdots,g_{k}\in O^{+}(NS(X))$ such that $x\in g_{i}\overline{P}$ and $B(x,r)\subset\bigcup_{i=1}^{k}g_{i}\overline{P}.$ In particular, we have
$B^{n-1}\cong B(x,r)\cap S\subset\bigcup\big{\{}g_{i}T:g_{i}\in\operatorname{Aut}(A(X)),T{\text{ is a side of }}\overline{P},{\text{and }}x\in gT\big{\}}.$
Hence there exists a side $T$ of $g\overline{P}$ such that $T^{\circ}\cap S\neq\varnothing.$ By [18, Theorem 6.3.8], we have $T\subset S.$ By $x\in T\cap\Sigma$, then $T$ is a vertical side of $g\overline{P}.$ Therefore, we have $\infty\in T\subset S$. ∎
There is a natural one-to-one correspondence from the set of maximal parabolic subgroups of $\operatorname{Aut}_{s}(X)$ to the rational ray of $\partial A(X)$. By multiplying a positive integer, we have a primitive element in $NS(X)$. From this argument and Lemma 5.14, we have the following proposition:
Proposition 5.17.
Let $P$ be a maximal parabolic subgroup of $\operatorname{Aut}_{s}(X).$ Then there exists an element $w\in W(X)$ such that the fixed point of $wPw^{-1}$ is an elliptic divisor $e$. Moreover, the rank of $P$ is equal to the rank of the Mordell-Weil group of the elliptic fibration $f_{e}$ associated with $e$.
5.2. The blown-up boundary of ample cones of K3 surfaces
The notion of the blown-up coronae for relatively hyperbolic groups (a generalization of geometrically finite groups) is defined in the paper of Fukaya-Oguni [11], and Fukaya [10] established Bestvina-Mess type formula for relatively hyperbolic groups. We can immediately obtain Theorem 1 by applying Fukaya’s result [10, Corollary B].
For readability to algebraic geometers, in this section, we give a simplified version of Fukaya’s result, using the geometry of ample cones of K3 surfaces and hyperbolic geometry. Note that the proof method and some definitions in this section are not itself new.
Definition 5.18.
Let $U$ be an open subset of $\mathbb{R}^{n}$, and let $p$ be a point in $U$. We denote the $(n-1)$-dimensional unit sphere in $\mathbb{R}^{n}$ by $S^{n-1}$. We put
$\operatorname{Bl}_{p}(U)=\{(x,t)\in U\times S^{n-1}:|x-p|t=(x-p)\},$
where ||$\cdot$|| ⋅ | is the Eulidean norm. We define a function $\pi_{p}:\operatorname{Bl}_{p}(U)\rightarrow U$ by setting $\pi_{p}(x,t)=x.$ We call the function $\pi_{p}$ the blowing-up of $U$ at a point $p$ (to $S^{n-1}$). For a subset $A$ of $U$, we call the closure of $\pi_{p}^{-1}(A\backslash\{p\})$ in $\mathbb{R}^{n}$ the strict transform of $A$ at a point $p$, denoted by $\operatorname{St}_{p}(A)$.
Since $\pi_{p}$ restricts to a homeomorphism from $\mathrm{Bl}_{p}(U)\backslash\pi^{-1}_{p}(p)$ to $U\backslash\{p\},$ we identify these two spaces.
Definition 5.19.
Let $\mathcal{P}=\{p_{1},\dots,p_{m}\}$ be a subset of $U$. For each $p\in\mathcal{P}$, we take an open neighborhood $U_{p}$ of $p$ such that $U_{p}\cap\mathcal{P}=\{x\}.$ Let $\pi_{p_{i}}:\operatorname{Bl}_{p_{i}}(U_{p_{i}})\rightarrow U_{p_{i}}$ be the blown-up of $U_{p_{i}}$ at a point $p_{i}$.
We define $\mathrm{Bl}_{\mathcal{P}}(U)$ by
$\mathrm{Bl}_{\mathcal{P}}(U)=U\backslash\mathcal{P}\cup({\bigcup_{p\in\mathcal{P}}\mathrm{Bl}_{p}(U_{p})}).$
We define the function $\pi_{\mathcal{P}}:\operatorname{Bl}_{\mathcal{P}}(U)\rightarrow U$ by setting
$\pi_{\mathcal{P}}(x)=\left\{\begin{array}[]{ll}x&(x\in U\backslash\mathcal{P})\\ \pi_{p}(x)&(x\in\operatorname{Bl}_{p}(U_{p}),p\in\mathcal{P})\end{array}\right.$
We call the function $\pi_{\mathcal{P}}$ the blowing-up of $U$ at a subset $\mathcal{P}$. For a subset $A$ of $U$, we call the closure of $\pi_{\mathcal{P}}^{-1}(A\backslash\mathcal{P})$ in $\mathbb{R}^{n}$ the strict transform of $A$ at a subset $\mathcal{P}$, denoted by $\operatorname{St}_{\mathcal{P}}(A)$.
Definition 5.20.
Let $\Gamma$ be a geometrically finite subgroup of $M(B^{n})$, and let $C$ be the set of cusped limit points of $\Gamma$. For finite subsets $\mathcal{P}$ and $\mathcal{Q}$ of $C$ with $\mathcal{P}\subset\mathcal{Q}$, there exists the narural projection $\pi_{\mathcal{P},\mathcal{Q}}:\operatorname{Bl}_{\mathcal{Q}}(\mathbb{R}^{n})\rightarrow\operatorname{Bl}_{\mathcal{P}}(\mathbb{R}^{n})$ such that $\pi_{\mathcal{P}}\circ\pi_{\mathcal{P},\mathcal{Q}}=\pi_{\mathcal{Q}}.$
[…]
Lemma 5.21.
The action of a gemetrically finite subgroup $\Gamma$ of $M(B^{n})$ on $\mathbb{R}^{n}$ (resp. $B^{n}$, $\partial B^{n}$) is extended to an action on $\operatorname{Bl}_{C}(\mathbb{R}^{n})$ (resp. $\overline{B}^{bl}$, $\partial B^{bl}$).
Proof.
Since $\Gamma$ preserves the cusped limit point of $\Gamma$, the action of $\Gamma$ on $\mathbb{R}^{n}$ can be extend to that on $\operatorname{Bl}_{C}(\mathbb{R}^{n}).$ As $\Gamma$ preserves $\overline{B^{n}}$ and $\partial B^{n}$, $\Gamma$ acts on $\overline{B}^{bl}$ (resp. $\partial B^{bl}$) ∎
Definition 5.22.
Let $\Gamma$ be a geometrically finite subgroup of $M(B^{n})$. A point $a$ of $\partial B^{bl}$ is a limit point of $\Gamma$ in $\partial B^{bl}$ if there is a point $x\in B^{n}$ and there exists a sequence $\{g_{i}\}_{i\in\mathbb{N}}$ of $\Gamma$ such that $a=\lim g_{i}x$. The set of limit points of $\Gamma$ is called the blown-up limit set of $\Gamma$, denoted by $\Lambda^{bl}(\Gamma).$
Lemma 5.23.
Let $\eta:U^{n}\rightarrow B^{n}$ be the standard transformation from $U^{n}$ to $B^{n}$. For non-zero vector $v\in\partial U=E^{n-1}$ and $a\in\overline{U^{n}}$, we define a half line $l:[0,\infty)\rightarrow\overline{U^{n}}$ by $l(t)=tv+a$, and put $L(t)=(\eta\circ l)(t).$ Then
$\operatorname{St}_{e_{n}}(L(t))\cap\pi^{-1}(e_{n})=\Bigl{\{}(e_{n},\frac{v}{|v|})\Bigr{\}}\subset\operatorname{Bl}_{e_{n}}(\mathbb{R}^{n}).$
In particular, the intersection of $\operatorname{St}_{e_{n}}(L(t))$ with $\pi^{-1}(e_{n})$ is determined by the direction of $v$.
Proof.
By direct calculation, we obtain
$L(t)-e_{n}=\frac{2(tv+\rho(a)-e_{n})}{|tv+\rho(a)-e_{n}|^{2}}.$
Then
$\lim_{t\rightarrow\infty}\frac{L(t)-e_{n}}{|L(t)-e_{n}|}=\lim_{t\rightarrow\infty}\frac{tv+\rho(a)-e_{n}}{|tv+\rho(a)-e_{n}|}=\frac{v}{|v|}.$
Hence, $\operatorname{St}_{e_{n}}(L)=L\cup\{(e_{n},\frac{v}{|v|})\}.$
Corollary 5.24.
Let $V$ be a vector subspace of $\mathbb{R}^{n}$, and let $a$ be a point of $\overline{U^{n}}$. Let $A$ be an open $r$-neighborhood of $V+a$ in $U^{n}$ for a positive number $r$. Then,
$\operatorname{St}_{e_{n}}(\eta(A))\cap\pi^{-1}_{e_{n}}(e_{n})=\Bigl{\{}(e_{n},\frac{v}{|v|})\in\operatorname{Bl}_{e_{n}}(\mathbb{R}^{n}):v\in V\backslash\{0\}\Bigr{\}}\cong S^{r-1},$
where $r$ is the dimension of the vector space $V$.
Lemma 5.25.
Let $\Gamma$ be a geometrically finite subgroup of $M(B^{n})$. For every cusped limit point $p$ of $\Gamma$, we have
$\Lambda^{bl}(\Gamma)\cap\pi_{C}^{-1}(p)=\Lambda^{bl}(\Gamma_{p})\cong S^{r-1},$
where $r$ is the rank of the stabilizer group $\Gamma_{p}$, i.e., the rank of maximal abelian subgroup of $\Gamma_{p}$.
Proof.
By conjugating $\Gamma$ in $M(B^{n})$, we can suppose $p=e_{n}.$ Let $\eta:U^{n}\rightarrow B^{n}$ be the standard transformation. The group $\Gamma$ acts on $U^{n}$ by setting $\gamma(x)=(\eta^{-1}\gamma\eta)(x)$ for any $\gamma\in\Gamma.$
First, we show that $\Lambda^{bl}(\Gamma_{p})\cong S^{r-1}.$ There is a $\Gamma_{\infty}$-inveriant $r$-plane $Q$ of $E^{n-1}$ such that $Q/\Gamma_{\infty}$ is compact. For any $r>0$, the $r$-neighborhood $N(Q,r)$ of $Q$ in $U^{n}$ is $\Gamma_{\infty}$-invariant. The $r$-plane $Q$ has a description of the form $Q=V+a,$ where $V$ is the $r$-dimensional vector subspace of $E^{n-1}$ and $a$ is a vector in $Q$. By Corollary 5.24, we have
$\Lambda^{bl}(\Gamma_{p})\subset\{(e_{n},\frac{v}{|v|})\in\operatorname{Bl}_{e_{n}}(\mathbb{R}^{n}:v\in V\backslash\{0\}\}.$
Conversely, for $v\in V\backslash\{0\}$ and any integer $n>0$, there is an element $\gamma_{n}$ of $\Gamma_{\infty}$ such that $\gamma_{n}a\in N(nv+a,R),$ where $R$ is the Euclidean diameter of a fundamental domain of $Q$ for $\Gamma_{\infty}.$ By Lemma 5.23,
$\lim_{n\rightarrow\infty}\gamma_{n}x=\frac{v}{|v|}\in\Lambda^{bl}(\Gamma_{p}).$
Hence, we have
$\Lambda^{bl}(\Gamma_{p})=\{(e_{n},\frac{v}{|v|})\in\operatorname{Bl}_{e_{n}}(\mathbb{R}^{n}):v\in V\backslash\{0\}\}\cong S^{r-1}$
Secondly, we show that $\Lambda^{bl}(\Gamma)\cap\pi_{C}^{-1}(p)=\Lambda^{bl}(\Gamma_{p}).$ Let $a$ be a point in $\Lambda^{bl}(\Gamma)\cap\pi_{C}^{-1}(p).$ There exist a point $x\in B^{n}$ and a sequence $\{g_{i}\}$ of $\Gamma$ such that $a=\lim_{n\rightarrow\infty}g_{i}x.$ If the sequence $\{g_{i}\}$ has infinitely many elements of $\Gamma_{p}$, by the definition of $\Lambda^{bl}(\Gamma_{p})$, we have $a\in\Lambda^{bl}(\Gamma_{p}).$ Otherwise, there exists a proper cusped region $U(Q,r)$ such that $g_{i}x\in N(Q,r)$ for infinitely many $i\in\mathbb{N}.$ By Corollary 5.24, $a$ is in $\Lambda^{bl}(\Gamma_{p}).$ Hence we have $\Lambda^{bl}(\Gamma)\cap\pi_{C}^{-1}(p)\subset\Lambda^{bl}(\Gamma_{p}).$ The converse inclusion is obvious. ∎
Lemma 5.26.
For any $x\in B^{n}$,
$\Lambda^{bl}(\Gamma)=\overline{\Gamma x}\cap\partial B^{bl}.$
In particular, the blown-up limit set of $\Gamma$ is a $\Gamma$-inveriant closed subset of $B^{bl}$.
Proof.
By the definition of $\Lambda^{bl}(\Gamma)$, it is obvious that $\Lambda^{bl}(\Gamma)\supset\overline{\Gamma x}\cap\partial B^{bl}.$
Conversely, for any $a\in\Lambda^{bl}(\Gamma)$, we put $\overline{a}=\pi_{C}(a)\in\Lambda(\Gamma)$. If $a$ does not lie in the set $C$ of cusped limit points of $\Gamma$, we can show that $a$ is a limit point of $\Gamma$ in $\partial B^{bl}$ in the same way as a limit point of $\Gamma$ in $\partial B^{n-1}$. If $a\in C$, applying Lemma 5.25 to the cusped limit point $a$,
$a\in\Lambda^{bl}(\Gamma)\cap\pi^{-1}_{C}(\overline{a})=\Lambda^{bl}(\Gamma_{\overline{a}})=\overline{\Gamma_{\overline{a}}x}\cap\partial B^{bl}\subset\overline{\Gamma x}\cap\partial B^{bl}.$
Therefore $\Lambda^{bl}(\Gamma)\subset\overline{\Gamma x}\cap\partial B^{bl}.$
Lemma 5.27.
Let $\Gamma$ be a non-elementary geometrically finite subgroup of $M(B^{n})$. Then,
$\Lambda^{bl}(\Gamma)=\operatorname{St}_{C}(\Lambda(\Gamma)).$
Proof.
By Lemma 5.26, we have $\Lambda^{bl}(\Gamma)\supset\overline{\pi^{-1}_{C}(\Lambda(\Gamma)\backslash C)}=\operatorname{St}_{C}(\Lambda(\Gamma)).$
Conversely, it is enough to show that $\operatorname{St}_{C}(\Lambda(\Gamma))\cap\pi^{-1}_{C}(p)\subset\Lambda^{bl}(\Gamma_{p})$ for each $p\in C.$ By conjugating $\Gamma$ in $M(B^{n})$, we can suppose $p=e_{n}$. As the similar way in Lemma 5.25, we pass the upper half-space model $U^{n}$, and $\Gamma$ acts on $U^{n}$. Let $U(Q,r)$ be a proper cusped region based at $\infty$. Then $\Lambda(\Gamma)\subset N(Q,r)\cup\{\infty\}.$ There exist a vector subspace $V$ of $E^{n-1}$ and a vector $a$ of $Q$ such that $Q=V+a.$ By Corollary 5.24 and Lemma 5.25 we have
$\operatorname{St}_{C}(\Lambda(\Gamma))\cap\pi^{-1}_{C}(p)\subset\big{\{}(e_{n},\frac{v}{|v|}):v\in V\backslash\{0\}\big{\}}=\Lambda^{bl}(\Gamma_{p})$
Definition 5.28.
Let $X$ be a K3 surface, and let $V$ (resp. $C_{X}$) be the same as in Definition 5.11. We denote the strict transform of $\partial A_{X}$ (resp. $\overline{A_{X}}$) at $C_{X}$ by $\partial A_{X}^{bl}$ (resp. $\overline{A_{X}}^{bl}$).
The following result is given by Baragar [3]. By using the geometrically finiteness of $O^{+}(NS(X))$, we can give another proof of the result.
Theorem 5.29 ( [3, Theorem 3.3] ).
For a K3 surface $X$, the set $\partial A_{X}\backslash\Lambda_{X}$ is equal to $V\backslash C_{X}=\{c\in V:\operatorname{Aut}(A(X))_{c}{\text{ is finite}}\}.$
Proof.
Let $x$ be a point of $A_{X}.$ For any $y\in\partial A_{X}$, we denote the hyperbolic line connecting $x$ to $y$ by $l_{x,y}.$
Let $P$ be the Dirichlet domain of $O^{+}(NS(X))$ centered at $x$. Then there exists a horoball $U(c)$ based at $c$ such that
  1. (1)
    $\{\overline{U(c)}:c\in V_{NS}\}$ are mutually disjoint,
  2. (2)
    $gU(c)=U(gc)$ for any element $g\in O^{+}(NS(X))$,
  3. (3)
    $\{\overline{U(c)}\backslash{c}\}$ is locally finite in $\mathbb{H}^{n}_{X}$,
where the set $V_{NS}$ is the same as in Definition 5.11. We put $K=\overline{P}\backslash\bigcup_{c\in C}U(c).$ Then, by Lemma 5.10, $\overline{A_{X}}^{H}\backslash\bigcup_{c\in C}U(c)=\operatorname{Aut}(A(X))\cdot K.$
If $y\notin C$, since cusped regions $\{U(c):c\in V_{NS}\}$ are mutually disjoint, there exists a sequence $\{a_{n}\}_{n\in\mathbb{N}}$ and $\gamma_{n}\in\operatorname{Aut}(A(X))$ such that $a_{n}$ converges to infinity as $n\rightarrow\infty$ and $l_{x,y}(a_{n})\in\gamma_{n}K.$ Since the Euclidean diameter of $\gamma_{n}K$ goes to zero as $n$ goes to infinity, we have $y=\lim_{n\rightarrow\infty}l_{x,y}(a_{n})=\lim_{n\rightarrow\infty}\gamma_{n}x\in\Lambda_{X}.$
If $y\in C$ and the rank of the stabilizer group $\operatorname{Aut}(A(X))_{y}$ is 0, by Lemma 5.13, the point $y$ is isolated in $\partial A_{X}.$ In particular, $y$ is not in the limit set $\Lambda_{X}$.
Otherwise, $y$ is in $C_{X}$, where $C_{X}=\{c\in C:\operatorname{vcd}W(X)_{c}<n-1\}.$ Then $y$ is a bounded parabolic point. In particular, $y$ is in the limit set $\Lambda_{X}.$
Definition 5.30.
Let $X$ be a K3 surface. In Definition 5.12, we defined the limit set of $X$, denoted by $\Lambda_{X}.$ We denote the blown-up limit set of $\operatorname{Aut}_{s}(X)$ by $\Lambda_{X}^{bl}.$
Corollary 5.31.
If the symplectic automorphism group $\operatorname{Aut}_{s}(X)$ of a K3 surface $X$ is non-elementary, then $\partial A_{X}^{bl}\backslash\Lambda^{bl}_{X}=V\backslash C_{X}.$
Proof.
By Theorem 5.29, we have $\partial A_{X}\backslash V=\Lambda(\Gamma)\backslash C_{X}.$ Then, by Lemma 5.27, we have the desired equality. ∎
Let $C$ be the set of cusped limit points of $\operatorname{Aut}(A(X)).$ By Lemma 3.32 and Lemma 5.16, there exist cusped regions $U(c)$ based at $c\in C$ such that
  1. (1)
    $\{U(c):c\in C\}$ are mutually disjoint,
  2. (2)
    $gU(c)=U(gc)$ for any $g\in\operatorname{Aut}(A(X))$ and $c\in C,$
  3. (3)
    $\{\overline{U(c)}\backslash\{c\}\}$ is locally finite in $B^{n}\cup O(\operatorname{Aut}(A(X)))$.
  4. (4)
    $U(c)$ meets just the sides of $\overline{A_{X}}$ incident with $c$.
Since a cusped region $U(c)$ is homeomorphic to its strict transformation on $\overline{B}^{bl}$, we identify $U(c)$ with its strict transformation. To apply the Bestvina-Mess formula (Theorem 4.11) to automorphism groups of K3 surfaces, we construct a homotopy on $\overline{A_{X}}^{bl}\backslash\cup_{c\in C}U(c)$.
Lemma 5.32.
Let $a$ be a point of $A_{X}$, and let $K(a,t)=C(a,t)\cap\overline{A_{X}}\backslash\cup_{c\in C}U(c),$ where $C(a,t)$ is the closed $t$-ball centered at $a$. For every $x\in\overline{A_{X}}\backslash(\cup_{c\in C}U(c)\cup C),$ let $S_{x}$ be the smallest (horo)sphere (based) centered at $x$ that meets $K(a,t)$. Then $K(a,t)\cap S_{x}$ is a singleton.
Proof.
Suppose that there exists $y_{1},y_{2}$ in $K(a,t)\cap S_{x}$ with $y_{1}\neq y_{2}.$ Then the hyperbolic line $[y_{1},y_{2}]$ connecting $y_{1}$ to $y_{2}$ is contained in $C(a,t)\cap\overline{A_{X}}.$ For any $z\in(y_{1},y_{2})$, we have $d(x,z)<d(x,y_{1})=d(x,y_{2}).$ Then we have $z\in\cup_{c\in C}U(c).$ Since cusped regions $U(c)$ are mutually disjoint, there exists $c\in C$ such that $(y_{1},y_{2})\subset U(c)$ and $y_{1},y_{2}\in K_{t}\cap\partial U(c).$
We now pass to the upper half-space model $U^{n}$ and conjugate $\operatorname{Aut}(A(X))$ so that $c=\infty.$ We denote by $\Sigma$ (resp. $w_{1},w_{2}$) the horoball (resp. the point) on $U^{n}$ corresponding to $\partial U(c)$ (resp. $y_{1},y_{2}$). By Lemma 5.16, the Euclidean segment $l$ connecting $y_{1}$ to $y_{2}$ is contained in $K(a,t)\cap\Sigma.$ This contradicts the minimality of $S_{x}$. ∎
Definition 5.33.
Let $K(a,t)$ be the same as in Lemma 5.32. We put
$U=\cup_{c\in C}U(c)\cup C.$
For every $x\in\overline{A_{X}}\backslash U$ and $t\geq 0$, we define the function
$\rho_{a}:[0,\infty)\times\overline{A_{X}}\backslash U\rightarrow\overline{A_{X}}\backslash U$
as follows: if $x\in K(a,t)$, we define $\rho_{a}(x,t)=x,$ otherwise, we define $\rho_{a}(x,t)$ by the unique point of $K(a,t)\cap S_{x}$.
Proposition 5.34.
Let $\rho_{a}$ be the same as above. Then $\rho_{a}$ is continuous.
Proof.
If $\rho_{a}$ is not continuous, there exists a sequence $\{(x_{n},t_{n})\}$ such that $(x_{n},t_{n})$ converges to $(x,t)$ and $\rho_{a}(x_{n},t_{n})$ converges to $y\neq\rho_{a}(x,t).$
If $x$ is in $B^{n}$, passing to a subsequence of $\{(x_{n},t_{n})\}$, we can assume that each $x_{n}$ is in $B^{n}$. By the definition of $\rho_{a}(x_{n},t_{n})$ (resp. $\rho_{a}(x,t)$), there exists $r_{n}$ (resp. $r$) such that $S(x_{n},r_{n})\cap K(a,t_{n})=\{\rho_{a}(x_{n},t_{n})\}$ (resp. $S(x,r)\cap K(a,t)=\{\rho_{a}(x,t)\}$). Passing to a subsequence of $(x_{n},t_{n})$, we can suppose that $r_{n}$ converges to $r_{0}$. Since $d((x_{n},t_{n}),\rho_{a}(x_{n},t_{n}))=r_{n}$, we have $d(x,y)=r_{0}$. By the definition of $\rho_{a}(x,t)$, we obtain $r_{0}>r$ and
$r_{n}-d((x_{n},t_{n}),\rho_{a}(x,t))$ $\geq r_{n}-d((x_{n},t_{n}),(x,t))-d((x,t),\rho_{a}(x,t))$
$=(r_{n}-r_{0})+(r_{0}-r)-d((x_{n},t_{n}),(x,t)).$
Hence, for a sufficiently large $n$, we have $r_{n}>d((x_{n},t_{n}),\rho_{a}(x,t))$. This contradicts the definition of $r_{n}$.
[…]
$|p-y|=r,\hskip 5.69054pt|x-p|\leq r,\hskip 5.69054pt{\text{and }}r\leq p_{n}.$
Since $x$ is in $\partial B^{n}$, we have that $p=(x_{1},\cdots,x_{n-1},r)$ and $|x-p|=r$. Since $|p-y|=r$, this contradicts the uniqueness of $\rho_{a}(x,t)$. ∎
Proposition 5.35.
$\rho_{a}$ extends to $\overline{\rho_{a}}:[0,\infty)\times\overline{A_{X}}^{bl}\backslash\cup_{c\in C}U(c)\rightarrow\overline{A_{X}}^{bl}\backslash\cup_{c\in C}U(c)$.
Proof.
Let $x$ be a point in $\overline{A_{X}}^{bl}\backslash\cup_{c\in C}U(c)$ with $\pi_{C}(x)=c.$ We now pass to the upper half-space model $U^{n}$ so that $a=e_{n}$ and $c=\infty$ by an isometry $\phi:B^{n}\rightarrow U^{n}.$ Let $\Sigma$ be the horosphere $\{x\in U^{n}:x_{n}=r\}$ based at $\infty$ corresponding to $U(c).$
If $C(e_{n},t)\cap\overline{A_{X}}\cap U(c)=\varnothing,$ we define $\overline{\rho_{a}}(x,t)$ by the intersection $S(e_{n},t)$ with $\mathbb{R}_{>0}e_{n}.$ If $C(e_{n},t)\cap\overline{A_{X}}\cap U(c)\neq\varnothing,$ there exists the half-line $l_{x}$ on $\Sigma$ starting at $re_{n}$ such that $\operatorname{St}_{C}(\phi^{-1}\circ l)\cap\pi_{C}^{-1}(c)=\{x\}.$ For any $t\geq 0$, there exists a sufficiently large $R_{t}>0$ such that $\rho_{a}(s,x)$ is in $\Sigma$ for any $s\in(0,t]$ and any $x$ with $|x|>R_{t}$. Then, if we define $\overline{\rho_{a}}(t,x)$ by the intersection of $K(a,t)$ with $l_{x},$ this is the desired extension of $\rho_{a}.$
Theorem 5.36.
If $\operatorname{Aut}_{s}(X)$ is non-elementary, then $(\overline{A_{X}}^{bl}\backslash\cup_{c\in C}U(c),\partial A_{X}^{bl})$ is a $\mathcal{Z}$-structure for $\operatorname{Aut}(X).$ In particular,
$\operatorname{vcd}(\operatorname{Aut}(X))=\dim\partial A_{X}^{bl}+1.$
Proof.
We define the homotopy $H:[0,1]\times\overline{A_{X}}^{bl}\backslash\cup_{c\in C}U(c)\rightarrow\overline{A_{X}}^{bl}\backslash\cup_{c\in C}U(c)$ setting
$H(t,x)=\left\{\begin{array}[]{ll}(t,x)&(t=1)\\ \rho_{a}(\log{\frac{t+1}{t-1}},x)&(t\neq 1)\end{array}\right.$
By Proposition 5.35 and the definition of $\rho_{a}$, $H$ is continuous. Hence $\overline{A}_{X}^{bl}\backslash\cup_{c\in C}U(c)$ is an AR, and $\partial A_{X}^{bl}$ is a $\mathcal{Z}$-set of $\overline{A}_{X}^{bl}\backslash\cup_{c\in C}U(c)$. Then $(\overline{A}_{X}^{bl}\backslash\cup_{c\in C}U(c),\partial A_{X}^{bl})$ is a $\mathcal{Z}$-structure for $\operatorname{Aut}_{s}(X)$ (resp. $\operatorname{Aut}(X)$). By Theorem 4.11, we have the desired equality. ∎
Corollary 5.37.
If $\Lambda_{X}$ is homeomorphic to a Cantor set and $X$ has an elliptic fibration whose rank is larger than 00, then
$\operatorname{vcd}(\operatorname{Aut}(X))=\max\operatorname{MW}(f:X\rightarrow\mathbb{P}^{1}),$
where $f$ runs all the elliptic fibrations of $X$.
Proof.
For any $c\in C$, we have $\Lambda^{bl}_{X}\cap\pi_{C}^{-1}(c)\cong S^{r-1},$ where $r$ is the rank of $\operatorname{Aut}(A(X))_{c}$. By Proposition 5.17, $r$ is equal to the rank of the Mordell weil group corresponding to $c$. If $c\notin C$, since $\Lambda_{X}$ is totally disconnected, $\{c\}$ itself is an open neighborhood. Then the topological dimension of $\Lambda^{bl}_{X}$ is equal to $\max\operatorname{MW}(f:X\rightarrow\mathbb{P}^{1})$. ∎
Example 5.38.
Let $X$ be the K3 surface whose Néron-Severi lattice $NS(X)$ is given by
$NS(X)=\begin{pmatrix}2&4&1\\ 4&2&0\\ 1&0&-2\\ \end{pmatrix}.$
Baragar [4] studied the automorphism group of this K3 surface and the growth of orbits of curves. He showed that $\Lambda_{X}$ is a Cantor set and $\operatorname{Aut}(A(X))$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}*\mathbb{Z}/2\mathbb{Z}*\mathbb{Z}/2\mathbb{Z}.$
Although the automorphism group of $X$ is completely determined in [4] as above and then this is not a new example of computing the virtual cohomological dimensions, we can also apply Corollary 5.37 to this K3 surface $X$.
5.3. The ample cones of K3 surfaces and sphere packings
Let $X$ be a K3 surface, and let $\phi:\mathbb{H}_{X}^{n}\rightarrow B^{n}$ be an isometry from $\mathbb{H}_{X}^{n}$ to the conformal ball model $B^{n}$. The isometry $\phi$ induces the natural homeomorphism from $\partial\mathbb{H}_{X}^{n}$ to $\partial B^{n}.$ For a $(-2)$-curve $\delta$ and an ample class $a$ of $X$, we set
$H_{\delta}=\{x\in\mathbb{H}_{X}^{n}:d(x,a)=d(x,s_{\delta}a)\},$
$H_{\delta}^{+}=\{x\in\mathbb{H}_{X}^{n}:d(x,a)<d(x,s_{\delta}a)\}.$
The closure of $H_{\delta}$ (resp. $H_{\delta}^{+}$) in $\overline{\mathbb{H}_{X}^{n}}$ is denoted by $\overline{H_{\delta}}$ (resp. $\overline{H}_{\delta}^{+}$).
Definition 5.39.
We define the sphere associated with $\delta$ by $S_{\delta}=\phi(\overline{H_{\delta}})\cap\partial B^{n}$, which is isomorphic to the $(n-2)$-sphere. We define $B_{\delta}=\phi(\overline{H}_{\delta}^{+}\backslash\overline{H_{\delta}})\cap\partial B^{n}.$ We call $B_{\delta}$ the open ball associated with $\delta$.
By Proposition 5.8, we have $\partial A_{X}=\cap_{\delta\in\Delta^{+}(X)}\overline{B_{\delta}},$ where $\overline{B_{\delta}}$ is the closure of $B_{\delta}$ in $\partial B^{n}.$
Definition 5.40.
The boundary $\partial A_{X}$ associated with the ample cone of $X$ is a sphere packing if for any two different $(-2)$-curves $\delta_{1}$, $\delta_{2}$ of $X$, their associated spheres $S_{\delta_{1}}$ and $S_{\delta_{2}}$ are either disjoint or tangent to each other.
$\partial A_{X}$ is called a connected sphere packing if $\partial A_{X}$ is a sphere packing, and for any two different $(-2)$-curves $\delta$, $\delta^{{}^{\prime}}$ of $X$, there exist $(-2)$-curves $\delta_{1},\dots,\delta_{n}$ of $X$ such that $\delta=\delta_{1}$, $\delta^{{}^{\prime}}=\delta_{n}$, and two spheres $S_{\delta_{i}}$ and $S_{\delta_{i+1}}$ are tangent for $i=1,\dots,n-1.$
Theorem 5.41.
Let $X$ be an elliptic K3 surface. If $\partial A_{X}$ is a connected sphere packing, and every elliptic divisor of $X$ corresponds to the tangent point of some two spheres associated with (-2)-curves, then
$\operatorname{vcd}(\operatorname{Aut}(X))=\max\{\operatorname{rk}\operatorname{MW}(f)\}=\rho(X)-3,$
where $f$ runs all elliptic fibrations of $X$.
Proof.
As $\partial A_{X}$ is connected, we can sort the set of the open balls associated with (-2)-curves to $\{B_{i}\}_{i\in\mathbb{N}}$ such that for any $i\in\mathbb{N}$, the intersection of $B_{i}$ with $\cup_{j=1}^{i-1}B_{j}$ is not empty. Let $\mathcal{P}_{m}$ be the set $\{p:p\in B_{i}\cap B_{j}{\text{ for some two distinct }}i,j\leq m\}.$ We define the following sets:
  • $U_{m}=\bigcap_{i=1}^{m}B_{i}$,
  • $F_{m}=\bigcap_{i=1}^{m}\overline{B_{i}}$,
  • $F_{m}^{bl}=\operatorname{St}_{\mathcal{P}_{m}}(F_{m})=$ the strict transform of $F_{m}$ at $\mathcal{P}_{m}$
[…]
For each $m\in\mathbb{N}$, $F_{m}^{bl}$ is a topological manifold with its interior $U_{m},$ and then $F_{m}^{bl}$ and $U_{m}$ are homotopy equivalent. Hence we have
(5.4) $H^{*}(\partial A_{X}^{bl},\mathbb{Z})\cong\lim_{m\to\infty}H^{*}(F_{m}^{bl},\mathbb{Z})\cong\lim_{m\to\infty}H^{*}(U_{m},\mathbb{Z}).$
For each $m\in\mathbb{N}$ and a sufficiently small $\epsilon>0$, we define $V_{m+1}$ by the $\epsilon$-neighborhood of $\partial B^{n}\backslash B_{m+1}$ in $U_{m}.$ Then we can see as follows:
  • $V_{m+1}$ is contractible,
  • $U_{m+1}\cup V_{m+1}=U_{m},$
  • $U_{m+1}\cap V_{m+1}$ is homotopy equivalent to the $(n-2)$-sphere with finite $k$ punctures,
where $k$ is the number of the set $\{p:p{\text{ is a tangent point of }}B_{i}{\text{ and }}B_{m+1}{\text{ for some }}i\leq m\}.$ By Mayer-Vietoris exact sequence and the $(n-2)$-sphere with $k$ punctures is homotopy equivalent to the wedge sum of $(k-1)$-copies of $S^{n-3}$, we have the following exact sequence:
$\cdots\rightarrow H^{*}(U_{m})\rightarrow H^{*}(U_{m+1})\oplus H^{*}(V_{m+1})\rightarrow H^{*}(\vee_{i=1}^{k-1}S^{n-3})\rightarrow H^{*+1}(U_{m})\rightarrow\cdots.$
Therefore, we can inductively conclude that
$H^{*}(U_{m})=0\hskip 8.53581pt(*\geq n-2).$
In particular, by Theorem 5.36 and the equation 5.4, we have
$\operatorname{vcd}\operatorname{Aut}(X)$ $=\max\{m:H_{c}^{m}(A_{X},\mathbb{Z})\neq 0\}$
$=\max\{m:H^{m}(\partial A_{X}^{bl},\mathbb{Z})\neq 0\}+1$
$\leq n-2=\rho(X)-3,$
where $H^{*}_{c}$ is the cohomology with compact support . Lemma 5.14 implies that there exists an elliptic fibration $f$ such that $\operatorname{rk}\operatorname{MW}(f)=\rho(X)-3.$
Example 5.42.
The Apollonian circle packing is generated as follows: we begin with four mutually tangent circles. There is four curvlinear triangles. We can put the inscribed circle for each curvlinear triangle, so that the curvlinear triangle separated into new curvlinear triangles. Inductively, we put inscribed circles for each curvelinear triangle, and the closure of the union of the inscribed circles is the Apollonian circle packing, which is shown in the following Figure.
Figure 1. The Apollonian circle packing
In [2], Baragar constructed a K3 surface $Y_{2}$ with ample cone the Apollonian circle packing, whose Néron-Severi lattice is given by
$\begin{pmatrix}-2&2&2&4\\ 2&-2&2&4\\ 2&2&-2&0\\ 4&4&0&0\\ \end{pmatrix}.$
By Theorem 5.41, we have $\operatorname{vcd}\operatorname{Aut}(Y_{2})=1.$
Example 5.43.
The Apollonian sphere packing is the three dimensional analog of the Apollonian circle packing. We begin with five mutually tangent spheres, and in the space between any four of them, we put the inscribed sphere. As the same manner in the previous example, we can inductively put a sphere so that this sphere touch other four spheres. The closure of the union of the inscribed spheres is the Apollonian sphere packing. Baragar [2] also constructed a K3 surface $Y_{3}$ with ample cone the Apollonian sphere packing, whose Néron-Severi lattice is given by
$\begin{pmatrix}-2&2&2&2&4\\ 2&-2&2&2&4\\ 2&2&-2&2&4\\ 2&2&2&-2&0\\ 4&4&4&0&0\\ \end{pmatrix}.$
By Theorem 5.41, we have $\operatorname{vcd}\operatorname{Aut}(Y_{3})=2.$